Grade 11Physical Chemistry

two tonnes of iron ore containing 94% of Fe2O3 produces iron in pure state. Calculate the weight of iron.

anwesha
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
. Fe2O3 --------> 2Fe + 3/2O2
0.94/160 moles         x moles of Fe
thus, from unitary method,
x = 0.94/160*2 moles 
x = 0.001175 moles
weight of fe = 0.01175*56 = 0.658 gm.